i need help with evaluating functions. Can you help me with this problem

f(x)=1/2x+2 the x value is -6

Answer by longjonsilver(2297)

Please read my Lesson on this site on functions, it will help you.

Anyway, f(x) = (1/2)x + 2

this says...whatever the value is, here it is written as "x", in the f(x) bit, then what you do is multiply it by (1/2) and then add 2...the equation!

so f(-6) = (1/2)(-6) + 2 which becomes
f(-6) = -3 + 2
f(-6) = -1

jon.